Your great uncle died and left you $125,000. Assuming you can get a constant rate of return of 5%, how much will you have when you want to retire in 45 years if interest is compounded monthly?

How much of the interest is due to interest on interest (or due to compounding)
Answer #1

Future value is calculated as:-

=FV(rate,nper,pmt,pv)

=FV(5%/12,45*12,,-125000)

=1180436.09

Interest =1180436.09-125000 =1055436.09